. If sales in 2015 were $15 million and sales in 2018 were $21 million then what must be the rate of growth if sales are continuously compounded?
Answer #1

The formula for continuous compounding:

A = Pert

Where, A = Future value = $21 million

P = Present value = $15 million

r = rate of growth

t = time or years = 2018-2015 = 3 years

21 = 15er-3)

1.4 = elr.3)

log(1.4) = log(er)

0.1461 = 3r.log(e)

0.1461 log(e)

0.1461 3.log(e)

0.1461 1.3029 =

ir= 0.1121

Therefore r = 11.21%